OpenStudy (anonymous):

What is the x-value in the solution of the system? x-4y+7z=-24 4x-3z=14 3x-2y=0

OpenStudy (anonymous):

What you would have to do is substitution... I think its been awhile. Since the first equation has x, y and z. Use the other two equations to replace y and z with x. So for 4x-3z=14 get z by itself so you can substitute that in for the first equation for z. -3z=14-4x Divide by -3 to get z \[z=\frac{ 14-4x }{ -3 }\]Now do that for 3x-2y=0 but get y by itself fro the same reason as before. -2y=-3x Divide by -2 to get y by itself \[y=\frac{ 3 }{ 2 }x\] Now plug those two back into the first equation to get \[x-4(\frac{ 3 }{ 2 }x)+7(-\frac{ 14 }{ 3 }+\frac{ 4 }{ 3 }x)=-24\]Then solve for x or get x by itself. Hope this helps :)

OpenStudy (anonymous):

no sorry its through process of elimination

OpenStudy (anonymous):

Oh my bad sry its been awhile.